# Is poisoning a real threat to LLM alignment? Maybe more so than you think

### v2

@@ 4.2 Evaluation
We use two forms of evaluation in this work. 1. We use a clean reward model learned from the non-poisoned clean dataset using the Bradly Terry formulation to [4] of the reward function. This model is similar to the reward model used in PPO-based RLHF methods. We use this reward model’s response rating to evaluate the poisoned model’s harmfulness. Regarding backdoor attacks, we use the difference between rating for the poisoned response (prompt + trigger) and clean response (prompt) as the poison score. In the case of non-backdoor attacks, we consider the difference between the clean and poisoned model’s response as the poison score. Here the clean reward model is a Llama 2 7B based model. 2. We also use GPT4 to rate the responses between $1-5$ given the context of harmfulness. We follow the works of [36] to give a context of different types of harmfulness and ask GPT 4 to rate the responses. For further details about the evaluation, refer to Appendix C. In the GPT4-based evaluations, the poison score corresponds to the rating given by GPT4 to the response from a model. We find that the clean reward-based evaluation is consistent with the GPT4-based evaluation. We performed evaluations on a set of 200 prompts that were sampled from the test set.
@@ 4.3 Results
DPO score-based attacks: As opposed to the PPO as shown in the work of [5] where, selecting poison points based on the highest reward differential between chosen and rejected responses didn’t result in an increase in the efficiency of the poisoning, in the case of DPO selecting points based on the DPO score resulted in an extraordinary increase in the poisoning efficacy. Rather than needing 4-5% of the data to poison the model via the DPO score-based selection, we achieved a similar level of poisoning in even as small as 0.5% of data points as seen in 1. For further results refer to Appendix B.
@@ Appendix A Semantic Diversity based attack
Appendix B DPO Score based Attacks
@@
Appendix C GPT4 based evaluation

### v4

@@ 4.2 Evaluation
We use two forms of evaluation in this work. 1. We use a clean reward model learned from the non-poisoned clean dataset using the Bradly Terry formulation to [4] of the reward function. This model is similar to the reward model used in PPO-based RLHF methods. We use this reward model’s response rating to evaluate the poisoned model’s harmfulness. Regarding backdoor attacks, we use the difference between rating for the poisoned response (prompt + trigger) and clean response (prompt) as the poison score. In the case of non-backdoor attacks, we consider the difference between the clean and poisoned model’s response as the poison score. Here the clean reward model is a Llama 2 7B based model. 2. We also use GPT4 to rate the responses between $1-5$ given the context of harmfulness. We follow the works of [36] to give a context of different types of harmfulness and ask GPT 4 to rate the responses. For further details about the evaluation, refer to Appendix D. In the GPT4-based evaluations, the poison score corresponds to the rating given by GPT4 to the response from a model. We find that the clean reward-based evaluation is consistent with the GPT4-based evaluation. We performed evaluations on a set of 200 prompts that were sampled from the test set.
@@ 4.3 Results
DPO score-based attacks: As opposed to the PPO as shown in the work of [5] where, selecting poison points based on the highest reward differential between chosen and rejected responses didn’t result in an increase in the efficiency of the poisoning, in the case of DPO selecting points based on the DPO score resulted in an extraordinary increase in the poisoning efficacy. Rather than needing 4-5% of the data to poison the model via the DPO score-based selection, we achieved a similar level of poisoning in even as small as 0.5% of data points as seen in 1. For further results refer to Appendix C.
@@
5 Effect of Defense
Most of the proposed defenses in the literature against both universal backdoor and non backdoor attacks are in the domain of image classification. Out of these proposed defences we consider the class of data anomaly detection based defences where the goal is to find and remove presumptive poisoned candidate points from the training data. These methods in the vision literature has exploited the loss, last layer embedding and gradients to identify anomalies. In this section we analyze on how effectively does these concepts translate into universal attacks on large language models.
Figure 9: A. Correlation
Figure 9: B. Poison detected
Figure 9: Spectral Defense: Embedding with the highest singular value (A) didn’t result in a separation between the clean and poisoned data points. The lack of correlation resulted in the spectral defense method not being able to filter out the the poisonous data points. Here Mistral 7B model was used. In the figure we check the percentage of total poisonous data points that were in the top 10% of the filtered points based on the correlation with the highest singular value. Here we tried using the representation for last two layers of the network individually
Figure 10: A. Epoch 2, 5 Clusters
Figure 10: B. Epoch 3, 5 Clusters
Figure 10: B. Epoch 4, 5 Clusters
Figure 10: A. Epoch 2, 10 Clusters
Figure 10: B. Epoch 3, 10 Clusters
Figure 10: B. Epoch 4, 10 Clusters
Figure 10: Gradient based defense: Here we check the percentage of poison data points that end up in the clusters when we cluster the last layer gradients. Bars in the green corresponds to the percentage of data in each cluster when compared against the total data points while the urple and the red bars respectively denote the percentage of poisoned and clean data points in the respective clusters. Gradients clusters don’t reveal the poisoned data points in a significant manner. Here we used a Llama 7B model trained with 5% poison. Here note that the clusters with predominantly poisoned data are negligible in size (1-5 data points) and removing them wouldn’t make too much of a difference in the poisoning efficacy.
Spectral Methods: When it comes to anomaly detection in backdoor attacks spectral method [37] work on the observation that the poisoned data points gets sufficiently separated from the clean data points in the embedding level when it’s correlation with the top singular vector of the covariance matrix of the dataset (the last layer embedding of the all the data points) is considered. Due to this observation filtering the top $n\%$ data points according to the correlation can result in the removal of most of the poisons. This happens due to the fact that the backdoor signals gets boosted in the last layers as they becomes a strong indicator for misclassification for the network. We find that these type of separation does not happen in language models even when a universal backdoors are sufficiently installed ($5\%$ poison) as seen in Figure 9
Figure 11: Loss based data filtering: High loss based data filtering methods (removing the top 10% data points) failed to detect the poisoned data efficiently. In the meanwhile low loss based filtering even though managed to detect 60% of the dpo score based poisoned data it failed to detect the poisons efficiently in case of both the random and gradient projection based attacks. Here we used a mistral model that was trained with 5% poisoned dataset.
Gradient based defense: For poisoning in general another line of gradient based defense [38] exploits the observation that the effective poisons separates from the other data points in the gradient space during the earlier training epochs thus dropping these low density gradients clusters can minimize the efficacy of the poisoning attack. Here the last layer gradients are sufficient enough for the defense. In case of the language models we find that this observation doesn’t hold. As seen Figure 10 gradient clusters don’t tend to hold a significant portion of the poisoned data points.
High Loss Removal: One of the earlier versions of data anomaly detection’s work on that fact that [39, 40, 41, 27] if the percentage of poisoned samples are smaller then we can identify them via the removal of high loss data points. As seen in 11 we found that high loss point removal doesn’t detect most of the poisons in all three cases of random, dpo score based and gradient projection based poisoning. Meanwhile, low loss based filtering was able to remove a significant amount of the dpo score based influential poisons but both the random and gradient projection based poisons where able to evade the detection.
@@ Appendix A Semantic Diversity based attack
Appendix B DPO vs PPO with random poisoning
Figure 13: PPO (1 epoch) with 10% random poisoning
Figure 13: DPO (2 epoch) with 10% random poisoning
Figure 13: Comparison of both PPO and DPO trained with a beta of 0.1 and 1, 2 epochs respectively. Evaluations were done with clean reward. PPO’s reward function was trained for 1 epcoh and PPO was trained for an additional epoch where as in the case of DPO it was in total trained for two epochs in order to account for the implicit learning of reward. Both PPO and DPO showed similar performance when it came to random poisoning
Appendix C DPO Score based Attacks
@@
Appendix D GPT4 based evaluation